OpenStudy (anonymous):

Algebra 1 anybody?????

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

doing quadratics... \[\frac{ x+3 }{ 5^{2}+12x-9 }\]

OpenStudy (anonymous):

Try this site http://mathway.com/

OpenStudy (agent0smith):

\[\Large 5x^{2}+12x-9\] Factor this. The method i like to use is (it may seem confusing at first, sorry): start by multiplying the first coefficient (5) by -9, to get -45. Now you need two numbers that add up to the middle coefficient (12) and multiply to equal -45... (15)(-3) = -45, and 15 - 3 =12 So now write it like this with the 15 and -3, since it was 5x^2 just put 5x in both brackets for now: \[\Large (5x + 15) (5x-3)\] Now, since at the start we multiplied by 5, divide that by 5 (divide the first brackets by 5)... \\[\Large \frac{ (5x + 15) (5x-3) }{ 5} = (x+3)(5x-3)\] So now... \[\Large \frac{ x+3 }{ 5x^{2}+12x-9 } = \frac{ x+3 }{ (x+3)(5x-3) }\]

OpenStudy (anonymous):

sooooooooooo it would be (in the end) \[\frac{ 1 }{ 5x-3 }\] ???????? @agent0smith

OpenStudy (anonymous):

SOMEBODY????? oops caps

OpenStudy (anonymous):

Can we all back up a tad??? Is the divisor 5^2+2x-9 as written by brookester6, or is it 5x^2+2x-9 as worked by @agent0smith ?

OpenStudy (anonymous):

uuuuuuuuuuuuhhhhhhh it's my way cuz i got it from my paper O.o

OpenStudy (anonymous):

Drat...even *I* screwed it up!!! Lemme rephrase: Is the divisor 5^2+12x-9 as written by brookester6, or is it 5x^2+12x-9 as worked by @agent0smith ? And since we crossed postings, then I guess it's 5^2+12x-9 ???

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

If that's the case, then aren't we left with a divisor of: 25 +12x - 9 Which is 16+12x ?? Which makes the problem: (x+3) / (12x + 15) ???

OpenStudy (anonymous):

which is 1 over 5x-3 :D) i think i gots it :P lol

OpenStudy (anonymous):

DAMMIT!!!! I have got CLUMSYTHUMBS TONIGHT!!!! I meant: (x+3) / (12x + 16)

OpenStudy (anonymous):

mmhmmm which 3 reduces off of everything and leaves you with (1)/(5x+3) :D) CHEESE!!! ahaha i gots it now thanks a bunches!!

OpenStudy (agent0smith):

@qweqwe123123123123111 I assumed it was a typo and she left out the x... she said "doing quadratics" remove the x and it's no longer a quadratic...

OpenStudy (anonymous):

hah true dat

OpenStudy (agent0smith):

@brookester6 wrote: doing quadratics... \[\Large \frac{ x+3 }{ 5^{2}+12x-9 }\] pretty safe to assume there's an x missing in there.

OpenStudy (anonymous):

-"she said "doing quadratics...remove the x and it's no longer a quadratic..." Agreed... :-) So I asked... I'm still rather confused. My brain tonight isn't working much better than my thumbs... :(

OpenStudy (anonymous):

haah that's okay mine either i had 5 tests and a quiz at school today... boy am i gonna sleep goooooood tonight :D)

OpenStudy (agent0smith):

@brookester6 \[\Large \frac{ x+3 }{ (x+3)(5x-3) } = \frac{ 1 }{ 5x-3 }\] Is correct! :)

OpenStudy (anonymous):

Yayy!! :-)

OpenStudy (anonymous):

yes dats correct :D) i used the math site